OpenStudy (anonymous):

An exponential function is: Select one: a. A function that describes a straight line. b. A function of the form x = by, where b is a positive real number other than 1 and x is any real number. c. A function of the form f(x) = bx, where b is a positive real number other than 1 and x is any real number. d. None of the above. Incorrect

OpenStudy (anonymous):

@Adjax

OpenStudy (anonymous):

a simple hint an exponential function incorporates the use of e...and I want you to check out these linKs: http://mathonweb.com/help_ebook/html/expoapps.htm https://www.khanacademy.org/math/algebra2/exponential_and_logarithmic_func/exp_growth_decay/v/exponential-growth-functions.. and after reading them you can answer the question yorself

OpenStudy (anonymous):

OK a fast/express answer is Choice D if you are in hurry

OpenStudy (anonymous):

that's not the answer I put incorrect next to it

OpenStudy (anonymous):

Choice d is the correct answer by the above specification...if you don't believe read those two links

OpenStudy (anonymous):

it envolves a a exponential b^x but mostly nowadays e^x is taken as exp

OpenStudy (anonymous):

I already got it wrong but get a second chance on the quiz to get it correct

OpenStudy (anonymous):

what do you mean..my proposed solution is wrong??..eh

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so what's the correct amnswer they are showin'?

OpenStudy (anonymous):

it doesn't say

OpenStudy (anonymous):

@BSwan :Where I got wrong?

OpenStudy (anonymous):

well exp function is y=b^x lolo hihi i cant see any other choise but D , so @Adjax ur corret

OpenStudy (anonymous):

but it said it was wrong

OpenStudy (anonymous):

wait ! is c like this ? c. A function of the form \(f(x) = b^x\), where b is a positive real number other than 1 and x is any real number.

OpenStudy (anonymous):

-_- next time take a snap shot, so we could see mathmatics symbols instead of copy paste , ok ?

OpenStudy (anonymous):

and i think B would be like this ? b. A function of the form \(x = b^y\), where b is a positive real number other than 1 and x is any real number. confirm if im correct about B,C if yes then the answer is C

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

uh...It means all the incoorect things is due to wrong info(not wrong but pasted as wrong)

OpenStudy (anonymous):

So am I going with C then ?

OpenStudy (anonymous):

yeah :P